# The Double-Ellipsoid Geometry of CLIP

### v1

@@ Abstract
Contrastive Language-Image Pre-Training (CLIP) is highly instrumental in machine learning applications within a large variety of domains. We investigate the geometry of this embedding, which is still not well understood. We examine the raw unnormalized embedding and show that text and image reside on linearly separable ellipsoid shells, not centered at the origin. We explain the benefits of having this structure, allowing to better embed instances according to their uncertainty during contrastive training. Frequent concepts in the dataset yield more false negatives, inducing greater uncertainty. A new notion of conformity is introduced, which measures the average cosine similarity of an instance to any other instance within a representative data set. We show this measure can be accurately estimated by simply computing the cosine similarity to the modality mean vector. Furthermore, we find that CLIP’s modality gap optimizes the matching of the conformity distributions of image and text.
@@ 1 Introduction
Multi-modal approaches, particularly Contrastive Language-Image Pre-Training (CLIP) [44], have significantly advanced various computer vision tasks, enabling applications that were hard to imagine just a few years ago. These advancements include generating high-quality images from user prompts [45, 41]; performing open-vocabulary classification [18], segmentation [34, 57] and detection [54]; captioning [40, 7], and even semantic editing [29, 27]. Much of this success can be attributed to the key advancements made in multi-modal learning, with CLIP playing a key role. Following the remarkable success of vision-language models (VLMs), a wide range of adaptations has emerged for additional domains such as 3D [19, 5, 58], video [49, 37] and audio [53, 15].
Despite the practical advances, the structure of CLIP’s latent space is not fully understood. While properties such as alignment, uniformity, and the well-studied modality gap have been explored, they do not fully capture the geometry underlying this multi-modal space. A common approach is to investigate the geometry of the raw embedding after projection to the unit hypersphere. It was thus argued that CLIP features lie on a low aperture cone [35]. Since projection is an information reducing operation which “flattens” the data, we examine the raw embedding, before normalization to a unit norm. We show that CLIP’s raw latent space exhibits a double-ellipsoid geometry, one for text and one for images. Both ellipsoids are shifted from the origin (see Fig. 1). The offset from the origin explains the narrow cone effect, while their separability explains the modality gap phenomenon [35, 11, 47].
We perform a data-driven analysis of CLIP’s latent space using the MS COCO [36] validation set, without relying on any prior assumptions. We observe that both distributions of image and text follow the thin shell phenomenon [31, 32], typical of high dimensional data, in which most of the mass concentrates within some range from the mean.
We then examine the afforded advantages of this geometry. It is shown that shifting the mass from the origin allows to control the sharpness of the response in contrastive learning. A known phenomenon in unsupervised contrastive learning is the existence of false negatives [3, 33, 56]. That is, negative instances in the batch possessing similar concepts which should not be embedded far apart. False negatives can considerably degrade performance [4, 22]. This can be mitigated to some extent by reducing the sharpness of the response. Thus, as uncertainty grows the network produces a more “blurred” response, reducing the overall loss for the entire dataset. We refer to this as semantic blurring. Practically, this is done by embedding uncertain concepts closer (in terms of cosine distance) to the mean vector. Uncertainty grows as the concept is more frequent, with higher probability of false negatives in the batch. We thus predict that frequent concepts are better aligned to the mean vector of the ellipsoid. We validate this experimentally achieving excellent agreement with our hypothesis. Finally, we suggest a new way for image interpolation afforded by the deeper understanding of the geometry. Our contributions can be summarized as follows:
@@
    We investigate the raw CLIP embedding showing each modality resides on a separable ellipsoid shell not centered at the origin.
@@
    We analyze the advantages of such structure, allowing to control the sharpness of the response in contrastive learning.
@@
    False negative cases are shown to benefit mostly from this embedding. We demonstrate that the contrastive loss is optimized near CLIP’s ellipsoid offsets, for the MS-COCO dataset.
@@
    A new definition of concept conformity is proposed, quantifying the popularity of a concept. We find that conformity is correlated to (cosine) similarity to the mean vector with a remarkable Pearson correlation of 0.9998 (for both image and text on MS-COCO).
@@
    We reveal that the distribution of the conformity of image and text are different and are best matched in CLIP’s current positioning of the ellipsoids, leading to a possible explanation for the modality gap.
@@
    We demonstrate how conformity can be used to quantify and rank image and text generators.
  7. 7.
    Leveraging this geometry, we develop vertical SLERP (vSLERP), which extends interpolation capabilities in spontaneous, training-free image generation.
@@ 2 Related Work
Contrastive representation learning is a powerful learning scheme, where models are trained to associate positive pairs (e.g., different views of the same image [6]) closely in the embedding space while pushing negative pairs (e.g., different images) apart. This simple yet effective approach has led to significant advances across a wide range of applications, i.e. image classification [6, 17], natural language processing [14, 30], 3D analysis [1, 55] and more.
The latent space induced by contrastive learning has been widely explored [2, 24, 52, 51], often conceptualized as a normalized hypersphere [51, 35]. Alignment and uniformity [51] are key properties of the Normalized Temperature-scaled Cross-Entropy (NT-Xent) loss [6]. Optimizing alignment and uniformity was shown to be crucial for preserving rich semantic structures in the latent space, leading to improvements in downstream performance across multiple domains [11].
With the rise of cross-modal contrastive models, such as CLIP [44], which align images and text in a shared embedding space, new challenges in latent space geometry have emerged. A notable issue is the modality gap [35], where embeddings from different modalities, such as images and text, are separated in the shared latent space. Moreover, the narrow cone effect was observed [35, 47], where features occupy only a limited portion of the angular space.
One of the main challenges in multimodal contrastive learning is of obtaining high-quality pairs. Web-scale datasets may include mismatched positive pairs [9, 12, 38, 50] or mislabeled negative pairs that are actually positive, referred to as false negatives [3, 33, 56]. Numerous approaches have emerged to address this challenge, such as by identifying and introducing hard negative examples [4, 8, 46, 25]. Our observations are that false negatives appear to play a significant role in forming the geometry of CLIP’s latent space.
@@ 3.1 Notations
We investigate CLIP space induced by ViT-B/32 encoders of $n=512$ dimensions, $\mathcal{X}=R^{n}$. Let $\mathcal{X}_{i}\subset\mathcal{X}$ be the image subspace and $\mathcal{X}_{t}\subset\mathcal{X}$ be the text (captions) subspace. We will see that they are different and in fact linearly separable. Let $v\in\mathcal{X}$ be a vector in this space. We denote by $v_{i}\in\mathcal{X}_{i}$ vectors of images and by $v_{t}\in\mathcal{X}_{t}$ vectors of text. The symbol $\mathop{\operatorname{\mathbb{E}}}$ stands for the expected value. The respective modality mean of image and text are $m_{i}=\mathop{\operatorname{\mathbb{E}}}_{\begin{subarray}{c}v_{i}\in\mathcal{X}_{i}\end{subarray}}[v_{i}]$ and $m_{t}=\mathop{\operatorname{\mathbb{E}}}_{\begin{subarray}{c}v_{t}\in\mathcal{X}_{t}\end{subarray}}[v_{t}]$. Let $\tilde{v}$ be the vector after subtraction of the respective modality mean. That is, for images, $\tilde{v}_{i}=v_{i}-m_{i}\,:\,v_{i}\in\mathcal{X}_{i}$ and for text $\tilde{v}_{t}=v_{t}-m_{t}\,:\,v_{t}\in\mathcal{X}_{t}$.
Our statistical analysis and many experimental results are based on MS-COCO [36] validation set, a common standard image-text dataset.
@@ 3.2.1 Thin shell theory
There is an intensive research related to the thin shell phenomenon [26, 43, 32, 23, 31]. Definitions of log concave distributions and isotropic random vectors appear in the supplementary. Since isotropic random vectors have a unit second moment for any $x(k),\,k=1,..n$, we get that the expected value of the squared Euclidean norm is
@@
As shown for example in [43], $\operatorname{\mathbb{E}}[\|x\|^{2}]\approx\operatorname{\mathbb{E}}^{2}[\|x\|]$, the expected norm of $x$ can be approximated by
@@
Recent studies have shown this bound for $\alpha=4$ [32], $\alpha=2.23$ [23] and most recently for $\alpha=\frac{1}{2}$ [31]. See more details in the above papers and the references therein. Essentially, this means the mass of the distribution is concentrated around a shell of radius $\sqrt{n}$.
@@ 4 Geometric Analysis
We begin by examining the statistics of image and text in the CLIP embedding space $\mathcal{X}$. This part is completely data-driven without any prior assumptions related to the training process. We focus on the raw CLIP embedding, which is the output of the encoder before $L_{2}$ normalization, i.e. before projection onto the unit hypersphere. This projection loses important information. It basically “flattens” the original geometry artificially, in a manner which is hard to analyze. More details and statistical data are provided in the Supp.
Let us first examine the known modality gap phenomenon [35] in the raw embedding. In Fig. 2, normalized histograms are shown for features 93, 134 and 494 of the CLIP latent vector. We get a bimodal distribution where image and text are clearly not drawn from the same distribution. For feature 93, for instance, the KL-divergence between the distributions is $\approx 301$ (a value above 1 implies a considerable deviation between the distributions). It was previously shown in [48, 11] that image and text can be separated linearly. We find there are actually 9 features which serve as sort of “tags” for image and text. More formally, we can define the measure of separability of a feature $\ell$ by
@@
A plot of the features sorted by separability is shown in Fig. 3 (left). In Fig. 2 (right) it is shown that the modalities are linearly separable (with $100\%$ accuracy) using only two such tag features (93 and 134), based on a linear SVM classifier (decision boundary shown in blue). We can thus state the following property (which holds exactly at least for MS-COCO):
@@
Figure 4: Illustration of statistics of image and text features after subtracting the mean for each modality. Top, histograms of 10 first features of image (first row) and text (second row). Bottom, histograms of $\|\tilde{v}\|$ of images and of text. Although each feature separately has most of the volume near zero, the distribution of $\|\tilde{v}\|$ has no volume below some threshold, indicating a thin shell phenomenon, typical for high dimensions.
In Fig. 4, we show some statistics of the features of $\tilde{v}_{i}$ and $\tilde{v}_{t}$ (where the mean is subtracted). The first 10 features in each vector are shown for both modalities. The distribution appears smooth, unimodal, with peak around zero. The norm $\|\tilde{v}\|$, however, is distributed within a small range (thin shell) such that there is no mass near zero.
@@
Figure 5: Normalized histograms of the variance of features (two left plots) indicates a long tail of high variance, hence an ellipsoid rather than an hypersphere. Off diagonal dominance, Eq. 6, implies certain strong feature correlations, or a tilted ellipsoid.
@@
A linear transformation can be applied to the embedding to transform the ellipsoid into a sphere. This process is called whitening [10, 28]. Analysis of whitened CLIP is in the supplementary, providing indications that multivariate normal distribution is a rather crud approximation.
@@ 5 Loss behavior on a double-ellipsoid
In this section, we validate that a shifted double-ellipsoid structure achieves optimality in terms of the CLIP contrastive learning loss.
For a batch containing $M$ image-text pairs, we denote by $\bar{v}_{i}^{j}=\frac{v^{j}_{i}}{\|v^{j}_{i}\|}$ and $\bar{v}_{t}^{j}=\frac{v^{j}_{t}}{\|v^{j}_{t}\|}$ the normalized image and text features of the $j$-th pair in the batch respectively. The multi-modal learning loss used in CLIP is the normalized temperature-scaled cross entropy loss (NT-Xent), a variation of InfoNCE [42] loss:
@@
As observed by [51], the loss can be decomposed into two terms: (1) Alignment, which encourages high cosine similarity for positive pairs, and (2) Uniformity, which encourages low cosine similarity among negative samples.
@@
Figure 6: Loss with respect to position of center of embedding. The parameter $\alpha$ controls the center of embedding, see Eq. 9 ($\alpha=0$ is the current non-origin-centered position of CLIP). (Top). For non-origin-centered position, the unified loss is optimally balancing between uniformity and alignment. (Bottom). The loss is ascending when only misclassified instances are examined, and descending for well-classified. The overall accuracy, contains both, well balanced at $\alpha\approx 0$.
@@
The values of $v_{t}$ remain unchanged.
The results show that the loss for correctly classified samples decreases monotonically with the shift toward the origin (i.e. that for perfect alignment as assumed for example by [35], shifting to the origin would be preferable). Conversely, the loss for misclassified samples increases. The overall loss balances alignment and uniformity for both correctly and misclassified samples, reaching an optimal $\alpha$ near zero. This aligns with the current CLIP embedding, though some deviation is expected, as the MS-COCO validation set is only an approximation of the full training set.
For completeness, the supplementary material includes the same experiment with the text ellipsoid shifted instead of the image, showing consistent behavior. To conclude:
{mdframed}
[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = gray!20, roundcorner = 4pt, ] Property 6: CLIP’s contrastive loss is optimized for shifted ellipsoids, balancing alignment and uniformity for both correctly classified and misclassified instances.

### v2

@@ Abstract
Contrastive Language-Image Pre-Training (CLIP) is highly instrumental in machine learning applications within a large variety of domains. We investigate the geometry of this embedding, which is still not well understood, and show that text and image reside on linearly separable ellipsoid shells, not centered at the origin. We explain the benefits of having this structure, allowing to better embed instances according to their uncertainty during contrastive training. Frequent concepts in the dataset yield more false negatives, inducing greater uncertainty. A new notion of conformity is introduced, which measures the average cosine similarity of an instance to any other instance within a representative data set. We show this measure can be accurately estimated by simply computing the cosine similarity to the modality mean vector. Furthermore, we find that CLIP’s modality gap optimizes the matching of the conformity distributions of image and text.
Machine Learning, Vision Language Models, Modality Gap, ICML
@@ 1 Introduction
Multi-modal approaches, particularly Contrastive Language-Image Pre-Training (CLIP) (Radford et al., 2021), have revolutionized computer vision tasks, enabling applications such as high-quality image generation (Ramesh et al., 2022; Nichol et al., 2021), open-vocabulary classification (He et al., 2023), segmentation (Liang et al., 2023; Yu et al., 2024), detection (Wu et al., 2023), captioning (Mokady et al., 2021; Cho et al., 2022), and semantic editing (Kim et al., 2022; Kawar et al., 2023). Beyond images, CLIP’s success extends to 3D (Hegde et al., 2023; Chen et al., 2023; Zhang et al., 2022), video (Tang et al., 2021; Luo et al., 2022), and audio domains (Wu et al., 2022; Guzhov et al., 2022).
Despite these advances, the structure of CLIP’s latent space remains poorly understood. Existing studies focus on properties like alignment, uniformity, and the modality gap (Liang et al., 2022) but overlook the geometry underlying this multi-modal space. The L2-normalization phase, which is integral when performing cosine similarity, practically reducing the data to the unit hypersphere. Since normalization is an information-reducing process, understanding the primary embeddings prior to normalization can reveal deeper insights into the latent space geometry.
In this paper, we propose analyzing the pre-normalized CLIP primary embedding for three key reasons: (1) Enhancing downstream tasks. While $L^{2}$-normalization is integral to the cosine similarity used during training, the primary embedding is directly employed in critical downstream tasks, including image generation and semantic editing. Analysis of the latent geometry can enhance the performance of these tasks. (2) Semantic significance of magnitude. Despite the cosine similarity is agnostic to the norm, we observe that magnitude still plays a significant and meaningful role. Notably, the largest embeddings in MS-COCO correspond to unusual or exotic captions (e.g., “I am not sure what this image is”, see full histogram and examples in Figure 14 in the Appendix). (3) Deeper understanding of contrastive learning. CLIP is an exceptional semantic encoder achieved through a rather generic contrastive loss and huge training data. Investigating the solutions found by CLIP allows deeper insights on contrastive learning, possible approaches to tackle false negatives and may shed light on unresolved phenomena, such as the modality gap and the narrow cone effect (Liang et al., 2022).
Our analysis reveals that CLIP’s primary latent space exhibits a double-ellipsoid geometry, with one ellipsoid for images and another for text. Both are shifted from the origin (see Fig. 1), in line with the narrow cone effect and the modality gap (Liang et al., 2022; Fahim et al., 2024; Schrodi et al., 2024). Using the MS-COCO validation set (Lin et al., 2014), we show that both modalities exhibit the thin-shell phenomenon (Klartag, 2023; Klartag & Lehec, 2022), where most of the mass concentrates within a specific range from the mean.
This geometry affords several advantages. The offset from the origin allows CLIP to control the sharpness of its response in contrastive learning, mitigating false negatives (Byun et al., 2022; Li et al., 2022; Yang et al., 2022);instances that are conceptually similar but incorrectly treated as negatives. Frequent concepts with higher uncertainty are embedded closer to the mean vector, a phenomenon we term semantic blurring, reducing loss and improving performance. Our experiments confirm that frequent concepts are better aligned to the mean vector of the ellipsoid, achieving excellent agreement with our hypothesis.
Leveraging this deeper understanding, we introduce a new definition of concept conformity, quantifying how close a sample resides with respect to all others. We find a remarkable correlation between conformity and cosine similarity to the mean vector (Pearson correlation: 0.9998 for MS-COCO). Furthermore, we show that the distribution of conformity differs between modalities, with CLIP’s ellipsoid alignment offering a plausible explanation for the modality gap.
Our contributions are as follows:
@@
    We reveal that CLIP embeddings form separable ellipsoid shells for each modality, shifted from the origin.
@@
    We analyze the benefits of this structure, including its role in controlling sharpness in contrastive learning.
@@
    We show that frequent concepts benefit most from this geometry, optimizing the contrastive loss near the ellipsoid offsets for MS-COCO.
@@
    We define concept conformity and demonstrate its strong correlation with similarity to the mean vector, offering insights into semantic organization.
@@
    We highlight the role of conformity in explaining the modality gap and propose its use in ranking text and image generators.
@@
    We introduce vertical SLERP (vSLERP), an interpolation method leveraging the geometry of CLIP’s latent space.
@@ 2 Related Work
Contrastive representation learning is a powerful learning scheme, where models are trained to associate positive pairs (e.g., different views of the same image (Chen et al., 2020)) closely in the embedding space while pushing negative pairs (e.g., different images) apart. This simple yet effective approach has led to significant advances across a wide range of applications, i.e. image classification (Chen et al., 2020; He et al., 2020), natural language processing (Gao et al., 2021; Kim et al., 2021), 3D analysis (Afham et al., 2022; Xie et al., 2020) and more.
The latent space induced by contrastive learning has been widely explored (Arora et al., 2019; Ji et al., 2023; Wang et al., 2022; Wang & Isola, 2020), often conceptualized as a normalized hypersphere (Wang & Isola, 2020; Liang et al., 2022). Alignment and uniformity (Wang & Isola, 2020) are key properties of the Normalized Temperature-scaled Cross-Entropy (NT-Xent) loss (Chen et al., 2020). Optimizing alignment and uniformity was shown to be crucial for preserving rich semantic structures in the latent space, leading to improvements in downstream performance across multiple domains (Fahim et al., 2024).
With the rise of cross-modal contrastive models, such as CLIP (Radford et al., 2021), which align images and text in a shared embedding space, new challenges in latent space geometry have emerged. A notable issue is the modality gap (Liang et al., 2022), where embeddings from different modalities, such as images and text, are separated in the shared latent space. Moreover, the narrow cone effect was observed (Liang et al., 2022; Schrodi et al., 2024), where features occupy only a limited portion of the angular space.
One of the main challenges in multimodal contrastive learning is of obtaining high-quality pairs. Web-scale datasets may include mismatched positive pairs (Chun et al., 2022; Gadre et al., 2024; Maini et al., 2023; Wang et al., 2023) or mislabeled negative pairs that are actually positive, referred to as false negatives (Byun et al., 2022; Li et al., 2022; Yang et al., 2022). Numerous approaches have emerged to address this challenge, such as by identifying and introducing hard negative examples (Byun et al., 2024; Chuang et al., 2020; Robinson et al., 2020; Kalantidis et al., 2020). Our observations are that false negatives appear to play a significant role in forming the geometry of CLIP’s latent space.
@@ 3.1 Notations
We investigate CLIP space induced by ViT-B/32 encoders of $n=512$ dimensions, $\mathcal{X}=R^{n}$. Let $\mathcal{X}_{i}\subset\mathcal{X}$ be the image subspace and $\mathcal{X}_{t}\subset\mathcal{X}$ be the text (captions) subspace. We will reaffirm that they are different and in fact linearly separable (Schrodi et al., 2024; Liang et al., 2022). Let $v\in\mathcal{X}$ be a vector in this space. We denote by $v_{i}\in\mathcal{X}_{i}$ vectors of images and by $v_{t}\in\mathcal{X}_{t}$ vectors of text. The symbol $\mathop{\operatorname{\mathbb{E}}}$ stands for the expected value. The respective modality mean of image and text are $m_{i}=\mathop{\operatorname{\mathbb{E}}}_{\begin{subarray}{c}v_{i}\in\mathcal{X}_{i}\end{subarray}}[v_{i}]$ and $m_{t}=\mathop{\operatorname{\mathbb{E}}}_{\begin{subarray}{c}v_{t}\in\mathcal{X}_{t}\end{subarray}}[v_{t}]$. Let $\tilde{v}$ be the vector after subtraction of the respective modality mean. That is, for images, $\tilde{v}_{i}=v_{i}-m_{i}\,:\,v_{i}\in\mathcal{X}_{i}$ and for text $\tilde{v}_{t}=v_{t}-m_{t}\,:\,v_{t}\in\mathcal{X}_{t}$.
Our statistical analysis and many experimental results are based on MS-COCO (Lin et al., 2014) validation set, a common standard image-text dataset.
@@ 3.2.1 Thin shell theory
There is an intensive research related to the thin shell phenomenon (Kannan et al., 1995; Paouris, 2006; Klartag & Lehec, 2022; Jambulapati et al., 2022; Klartag, 2023). Definitions of log concave distributions and isotropic random vectors appear in the Appendix. Since isotropic random vectors have a unit second moment for any $x(k),\,k=1,..n$, we get that the expected value of the squared Euclidean norm is
@@
As shown for example in (Paouris, 2006), $\operatorname{\mathbb{E}}[\|x\|^{2}]\approx\operatorname{\mathbb{E}}^{2}[\|x\|]$, the expected norm of $x$ can be approximated by
@@
Recent studies have shown this bound for $\alpha=4$ (Klartag & Lehec, 2022), $\alpha=2.23$ (Jambulapati et al., 2022) and most recently for $\alpha=\frac{1}{2}$ (Klartag, 2023). See more details in the above papers and the references therein. Essentially, this means the mass of the distribution is concentrated around a shell of radius $\sqrt{n}$.
@@ 4 Geometric Analysis
We begin by examining the statistics of image and text in the CLIP embedding space $\mathcal{X}$. This part is completely data-driven without any prior assumptions related to the training process. We focus on the primary CLIP embedding, which is the output of the encoder before $L_{2}$ normalization, i.e. before projection onto the unit hypersphere. This projection loses important information. It basically “flattens” the original geometry artificially, in a manner which is hard to analyze. More details and statistical data are provided in the Appendix.
Figure 4: Statistics of image and text features after mean subtraction. Top: The first 10 features for image (top) and text (bottom). Bottom: Histograms of $\|\tilde{v}\|$ for images and text, showing a thin-shell phenomenon with no volume below a threshold, typical for high dimensions.
Let us first examine the known modality gap phenomenon (Liang et al., 2022) in the primary embedding. In Fig. 2, normalized histograms are shown for features 93, 134 and 494 of the CLIP latent vector. We get a bimodal distribution where image and text are clearly not drawn from the same distribution. For feature 93, for instance, the KL-divergence between the distributions is $\approx 301$ (a value above 1 implies a considerable deviation between the distributions). It was previously shown in (Shi et al., 2023; Fahim et al., 2024; Schrodi et al., 2024) that image and text can be separated linearly. We find there are actually 9 features which serve as sort of “tags” for image and text. More formally, we can define the measure of separability of a feature $\ell$ by
@@
A plot of the features sorted by separability is shown in Fig. 3 (left). Fig. 2 (right) shows that the modalities are linearly separable (with $100\%$ accuracy) using only two such tag features (93 and 134), based on a linear SVM classifier (decision boundary shown in blue). We can thus state the following property (which holds exactly for MS-COCO):
@@
In Fig. 4, we show some statistics of the features of $\tilde{v}_{i}$ and $\tilde{v}_{t}$ (where the mean is subtracted). To get impression, the first 10 features in each vector are shown for both modalities. The distribution appears smooth, unimodal, with peak around zero. The norm $\|\tilde{v}\|$, however, is distributed within a small range (thin shell) such that there is no mass near zero.
Figure 5: Normalized histograms of feature variance (left) show a long tail, indicating an ellipsoid rather than a hypersphere. Off-diagonal dominance (Eq. 6) suggests strong feature correlations, implying a tilted ellipsoid.
@@ 5 Loss behavior on a double-ellipsoid
In this section, we validate that a non-origin-centered double-ellipsoid structure achieves optimality in terms of the CLIP contrastive learning loss.
For a batch containing $M$ image-text pairs, we denote by $\bar{v}_{i}^{j}=\frac{v^{j}_{i}}{\|v^{j}_{i}\|}$ and $\bar{v}_{t}^{j}=\frac{v^{j}_{t}}{\|v^{j}_{t}\|}$ the normalized image and text features of the $j$-th pair in the batch respectively. The multi-modal learning loss used in CLIP is the normalized temperature-scaled cross entropy loss (NT-Xent), a variation of InfoNCE (Oord et al., 2018) loss:
@@
As observed by (Wang & Isola, 2020), the loss can be decomposed into two terms: (1) Alignment, which encourages high cosine similarity for positive pairs, and (2) Uniformity, encourages low cosine similarity among negative ones.
@@
Figure 6: Loss vs. embedding center position. The parameter $\alpha$ controls the embedding center (Eq. 9, with $\alpha=0$ as the current non-origin-centered CLIP position). (Top). The unified loss balances uniformity and alignment optimally for non-origin-centered positions. (Bottom). The loss increases for misclassified instances and decreases for well-classified ones, with balanced accuracy at $\alpha\approx 0$.
@@
The values of $v_{t}$ remain unchanged. Unlike the Embedding Shift Experiment in (Liang et al., 2022), here, the modalities are not directly shifted to each other, but to the origin.
The results show that the loss for correctly classified samples decreases monotonically with the shift toward the origin (i.e. that for perfect alignment as assumed for example by (Liang et al., 2022), shifting to the origin would be preferable). Conversely, the loss for misclassified samples increases. The overall loss balances alignment and uniformity for both correctly and misclassified samples, reaching an optimal $\alpha$ near zero. This aligns with the current CLIP embedding, though some deviation is expected, as the MS-COCO validation set is only an approximation of the full training set. For completeness, the Appendix includes the same experiment with the text ellipsoid shifted instead of the image, showing consistent behavior. To conclude: {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = gray!20, roundcorner = 4pt, ] Property 6: CLIP’s loss is optimized for non-origin-centered ellipsoids, balancing alignment and uniformity for both correct and misclassified instances.
Figure 7: Top: Example of segmentation score blur (right), common in semantic segmentation, as object-membership uncertainty increases. Bottom: Similarity histograms of normally distributed samples for the mean vector (blue) and the furthest vector from the mean (”extreme”, orange). Results are shown for a sphere centered near the origin (left) and one centered far from the origin (right). In contrastive learning, blur can be controlled by adjusting the sphere’s offset. Embedding vectors closer to the center induces blur, while positioning them farther away sharpens the response.